# A 2-step Framework for Automated Literary Translation Evaluation: Its Promises and Pitfalls

### v2

@@ 5.2 VERSE: Can LLMs Answer Literary Questions with Subtle Nuance?
Model Kendall Spearman MSE
gpt-4o .49 .54 0.61
claude-3-5-sonnet .51 .54 0.73
claude-3-5-haiku .37 .40 0.77
gpt-4o-mini .37 .40 0.61
claude-3-haiku .20 .21 0.82
gemma-2-27b-it .13 .14 0.71
gemma-2-9b-it .09 .09 0.71
Llama-3.1-8B-it .05 .05 1.08
COMET .23 .30
BERTScore .19 .24
BLEURT .19 .24
Table 5: (Step 2) VERSE: Performance comparison across different models on evaluating whether a particular criterion was satisfied in the translation or not. Usually, larger models do better than smaller ones. The results are for zero shot no CoT setting. Baselines are provided below.
As shown in Table 5, language models are much better at correlating with human preference in understanding literary translation than the traditional MT metrics, however, there remains a gap between expert human-level agreement (Table 2). In-depth analysis of GPT4-o (the best-performing model above) performance shows that the model has a particularly lower recall than inter-annotators for Score 1 than for Score 2 and 3 (Table 8), which means GPT4-o cannot effectively spot translations that critically fail to satisfy specific criteria. However, the recall drop was not as significant as in the case of honorifics in Step 1.
We identify the following key components of LLM’s ability to answer fine-grained questions:
  • CoT only helps certain models and hurts a few (Table 12).
  • Human written rubrics (Appendix D) are critical (Table 10) and correlation drops to near zero without them.
  • Providing a reference translation does not help evaluation, in fact not providing is slightly better (Table 13).
  • Providing few shot examples does not make a significant improvement across the models (Table 14).
Appendix H digs more into these findings.
@@ 6.1 Experiment Setting
For all the experiments in this section, we use the GPT4-o as the evaluator model for all experiments in zero-shot no CoT setting. We also did not use reference human translation because, in Step 2, reference-free configuration performs slightly better [^6] We also report the evaluation scores computed with reference translation in the prompt in Appendix H.
@@
To test the efficacy of our framework, we use several language models and Google Translate to translate 10 stories from our dataset. Following the best practices of Karpinska and Iyyer (2023), we prompt language models with five examples and then generate translations. We also append a GPT4o-generated summary of the story for context.
From here on, we use $\mathrm{LM}_{E}$ and $\mathrm{LM}_{G}$ to differentiate between the evaluator language model and translator language models.
6.2 Fine Grained Literary Evaluation
@@
VERSE (Step 2) of our framework permits a novel fine-grained interpretation into literary translation, which is otherwise hard to capture through a single number or traditional machine translation metric like fluency, adequacy, etc. Figure 3 shows one such use case of such approach to quickly inspect and visualize different dimensions present in the translation. Moreover, because the language model evaluates  10 questions per paragraph, VERSE estimations are also robust.
@@ 6.3 LLMs prefers LLM generated translation
Table 6 shows the aggregated scores for each Step 1 criteria for each model (see Table 16 for Step 2). Note that since our proposed framework can be run in reference free fashion, we also evaluate the given reference human translations with our framework.
The results are surprising. Overall, Claude-3.5 Sonnet, as $\mathrm{LM}_{G}$, gets the highest scores, followed by GPT4-o, with almost perfect scores in several dimensions. Human translation receives lower scores than the bigger models. To better understand what is going on, we sample 50 paragraphs from the story “What is Remembered", whose human translation receive a particularly lower score than the best performing $\mathrm{LM}_{G}$ translation by ($\mathrm{LM}_{G}$) Claude Sonnet 3.5 (See Figure 8). In a controlled A/B setting, one human annotator picked the gold translation as the better one 39 out of 50 times, and the Claude Sonnet 3.5 translation was picked only 4 times.
An inherent characteristic of human translation may explain why $\mathrm{LM}_{E}$ gives human translation lower scores. Human translations often deviate significantly from the surface-level linguistic features of their original text in order to properly convey the essence of the text. Human inspection reveals that this is indeed the case for the story examined above. In the cases of other stories, too, human translations often prove to read more naturally than the best machine translations in most cases, even when the best machine translations were sufficiently accurate.
@@
This suggests that while LLMs may be good at spotting significant differences in translation qualities, they still often fail to distinguish translations where the difference is more subtle and nuanced. This can be observed from the relatively lower scores provided to Google Translate and GPT3.5, which usually produce very poor Korean translations.
Transl. Model Hon. Syn. Lex. Con.
Human 98.97 88.21 89.74 83.08
GPT-4o 99.49 97.44 94.36 97.95
Claude-3.5-Sonnet 100.00 98.97 98.97 98.97
Claude-3.5-Haiku 100.00 94.87 96.41 96.92
Claude-3-Haiku 98.46 88.72 85.64 89.23
GPT-4-Mini 99.49 90.26 88.21 92.31
GPT-3.5 91.79 57.44 56.92 60.00
Google Translate 97.95 71.79 69.23 76.41
Table 6: Average scores (in percentage) across different models on four criteria of translation quality, computed over the 10 evaluation stories (501 paragraphs). Note that while LLMs can discriminate between very good and very bad translations (eg, GPT4o vs GPT3.5), its ability to distinguish between high-quality translations is limited.
@@
While we keep our analysis limited to English to Korean, the overall structure of evaluation can be used for other languages, provided LLMs have a reasonable understanding of the target language. Our framework is simple, easy to implement, and provides interpretability suited specifically for literary translation. Much of the pipeline, especially, VERSE is language agnostic and works without much human supervision. For RULER, we created the rubric with the Korean language in mind, however, RULER can easily be adopted with modification for other languages. While gaps remain in current LLMs’ ability to detect some critical issues in literary translation, we believe this paradigm will be helpful for future researchers and practitioners to evaluate literary translation.
@@ 7 Conclusion
Our work only tests the framework on one pair of languages, English to Korean. Future works should incorporate other language pairs, which provide different levels of performance. Especially, it remains to be seen how well LLMs can generate literary-specific questions for other non-English languages and then answer them. It is also important to see whether LLMs can take into account different cultural contexts of the source language in the translation. It is possible that LLMs will not be able to evaluate high-quality questions in low-resource languages or languages that they haven’t been trained on extensively. Similarly, it is also possible that the framework works even better in higher-resource languages.
Appendix A Datast detail
Our selected collection of short stories spans nearly two centuries, offering a survey of diverse literary movements and genres in the English language. From classic nineteenth-century gothic and popular suspense stories like Edgar Allan Poe’s “The Masque of the Red Death” (1842) and Charles Dickens’s “The Signal-Man” (1866), to modernist text like Virginia Woolf’s “The Legacy” (1944) and contemporary realist works such as Jhumpa Lahiri’s “Hell-Heaven” (2004) and Andrew Porter’s “Austin” (2017). The genre of speculative fiction tradition is represented by Ursula K. Le Guin’s “The Ones Who Walk Away from Omelas” (1973), Ted Chiang’s “Tower of Babylon” (1990), and Octavia Butler’s “Amnesty” (2003), all of which differ in writing style, while works like Charles Chesnutt’s “The Passing of Grandison” (1899), Katherine Mansfield’s “The Garden Party”(1922) and Tobias Wolff’s "The Deposition" (2006) represent works in which historical and cultural detail contribute greatly to their style. As a collection, these stories cover a wide-range of evolving narrative styles and techniques in English literature.
Paragraphs from five of the fifteen stories (shaded dark gray on top) were used to create the few-shot examples for all the experiments in Section 4 and 6. The other five works (lighter gray) were part of the human evaluation in 4 but their paragraphs were not part of the few-shot examples. We leave the five other stories solely for the purpose of our analysis in Section 6.
The names of the stories and other details can be found in Table 7.
Title Author #P Translator(s) Orig. Trans.
Austin Andrew J. Porter 26 Eun Young Min 2017 2024
The Passing of Grandison Charles W. Chesnutt 56 Ki Wook Han 1899 2010
The Ones Who Walk Away
@@ Appendix A Datast detail
The Garden Party Katherine Mansfield 44 Young Hee Kim 1922 2010
Real Artists Ken Liu 84 Seong Ju Jang 2011 2024
The Philosopher Sherwood Anderson 29 Seon H. Kim, Young W. Park 1919 16, 19
A Retrieved Reformation O. Henry 36 W. D. Kim, Jeong A Ko 1903 03, 14
Amnesty Octavia Butler 84 Soo Hyun Lee 2003 2016
Tower of Babylon Ted Chiang 107 Sang Hoon Kim 1990 2016
The Red Masque of Death Edgar Allan Poe 14 D. H. Jeon, S. H. Jeon 1842 09, 14
@@
Table 7: The details of the stories being used in our work. The number of paragraphs and when they were published are shown. Three stories in our corpus have two different translations, for which we included the names of both translators. Gray row stories were used for human evaluation experiments in Section 5.
Appendix B Details of Common Issues found in Machine Translation
Word/Phrase/Expression Level Issues
This type of error occurs most frequently and significantly impacts overall translation quality. While LLMs excel at literal translations, they struggle with natural word choices in the target language, particularly when dealing with figurative expressions. This often results in the selection of awkward or uncommon verbs in Korean; although the overall content remains comprehensible, these choices critically diminish the quality of translation. Notably, problems often arise with polysemous words or idiomatic expressions, seriously impeding reader comprehension. Additional common errors include inappropriate use of pronouns, such as referring to people as “{CJK*}UTF8mj그것 (it),” reflecting the essential difference in the linguistic structure between English and Korean.
Context Related Issues
This type of error presents more complex and serious challenges arising from cultural and historical differences between the source and target languages. For instance, because terms like “abolitionist” in “The Passing of Grandison” do not have an equivalent historical term in Korean language, LLMs end up offering incomprehensible expressions. Similarly, the expression “border states” often loses its geographical context (mistranslated as “{CJK*}UTF8mj국경주–states close to national borders” in Korean). LLMs also struggle with specialized terminology and historical references, as shown in mistranslations of terms like “bootjack” and “nightcap.”
Form-related issues
This third category includes subcategories concerning punctuation and grammar, as well as structure and formatting. Punctuation and grammar refer to problems with punctuation changes that alter meaning. For example, when LLM translation replaces semicolons with periods, it alters the original text’s meaning. Likewise, changes in the sentence structure in the translated text can lead to incomplete or awkward results. This subcategory includes problems with splitting long sentences unnecessarily, which results in awkward structures when handling complex sentences. Other issues in this category included structure and formatting. These appeared when formatting the dialogue and narrative flows. The incorrect handling of sentence structures in the LLM translation led to awkward sentences, negatively affecting readability. Form-related issues also include transliteration, consistency issues, and omissions. LLMs transliterate Western names inconsistently in the translation, resulting in different name variations throughout the translated text. (i.e., translating “Grandison” as “{CJK*}UTF8mj그라디슨" and “{CJK*}UTF8mj그란디손” within a single translation) LLM translations may eliminate a sentence and expression or add arbitrary content different from the original text. They also leave certain words untranslated, sometimes including various words or phrases in random languages within the translation.
Honorifics
Appendix C Detailed Rubric for Each Criteria (Step 1)
C.1 Lexical Choices
Proper use of idiomatic expressions and general word choices that demonstrate fluency in the target language
  • Score 5 The translated text chooses words in the target language that best suit the authorial intention of the original text, conveying meaning in a properly idiomatic and fluent manner. The translated text thereby provides readers with an equivalent reading experience and emotional impact when compared with the original text.
  • Score 4 The translated text mostly selects appropriate words that match the authorial intention, maintaining idiomatic expressions and fluency in the target language. The translated text generally preserves the reading experience and emotional effect of the original text.
  • Score 3 The translated text makes some less accurate or awkward word choices that compromise idiomatic fluency while demonstrating good comprehension of the original text. Lexical choices in the translation somewhat negatively affect readers’ experience and emotional engagement with the text.
  • Score 2 The translated text includes multiple inappropriate word choices that compromise idiomatic fluency to the extent that it makes it difficult to understand the authorial intention of the original text. Lexical choices in the translation somewhat negatively affect the reading experience and emotional engagement.
  • Score 1 Translated text makes clearly awkward word choices that reveal a critical lack of understanding of idiomatic and natural expressions in the target language, seriously hindering readers’ comprehension and engagement with the text. Lexical choices in the translation reflect critical miscomprehension of the authorial intent of the original text.
C.2 Proper Use of Honorifics in Dialogues
Selection of verb endings that accurately reflect the status of and relationships among the participants in a conversation, resulting in natural-sounding dialogue in the Korean language.
  • Score 5 The translated text properly captures the relationship between the speaker and the listener(s) of the conversation. The translated text also accurately reflects the culturally and historically specific elements such as titles and modes of appellations. The level of politeness aligns appropriately with the situational context, familiarity, and character.
  • Score 4 The translated text mostly captures the relationship between the speaker and the listener(s) of the conversation, with minor inaccuracies in culturally and historically specific elements such as titles and modes of appellations. The level of politeness mostly aligns with the situational context and social status of characters. There may be minor inconsistencies but the dialogue accurately represents the characters’ familiarity and relationships.
  • Score 3 The translated text somewhat captures the relationship between the speaker and the listener(s). But the translated text includes noticeable errors in culturally and historically specific elements such as titles and modes of appellations. The level of politeness partially aligns with the situational context and the social status of or relationships of the characters. There are some inconsistencies or awkwardness in the dialogue that are noticeable for native speakers of Korean.
  • Score 2 The translated text barely captures the relationship between the speaker and the listener(s). The translated text demonstrates significant misunderstandings of culturally and historically specific elements such as titles and modes of appellations. The level of politeness barely aligns with the situational context, leading to a misrepresentation of the characters’ relationships and awkwardness in the target language.
  • Score 1 The translated text completely misunderstands the relationship between the speaker and the listener(s) of the conversation and fails to reflect cultural and historical specificity of the context. The level of politeness in the used language fails to align with the situational context. Inconsistencies or misinterpretations of the appropriate degree of formality significantly undermine the dialogue’s naturalness in Korean.
C.3 Syntax and Grammar
Grammatical accuracy of sentences including proper translation of embedded clauses, conjunctions, sentence connectors, internal logic of sentences, and punctuations, which ultimately makes the translation read naturally in the target language.
  • Score 5 The translated text accurately conveys the meaning, logic, and nuances of the original text. Embedded clauses, conjunctions, and sentence connectors are translated meticulously so that the translation reads naturally in the target language without awkward or confusing sentence structures. The number of sentences in the translated text may vary from the number of sentences in the original text if it enhances the naturalness and readability of the translation.
  • Score 4 The translated text generally preserves the meaning, logic, and nuances of the original text. The translation of embedded clauses, conjunctions, and sentence connectors reads naturally, with only minor instances of awkwardness or ambiguity. Variations in the number of sentences mostly contribute to the naturalness and readability of the translation.
  • Score 3 The translated text partially preserves the meaning, logic, and nuances of the original text. Some embedded clauses, conjunctions, and sentence connectors are translated inaccurately, resulting in awkward sentence structures in the target language. The number of sentences in the translation varies from the number of sentences in the original text without a clear demonstration of the enhancement in readability and naturalness.
  • Score 2 The translated text falls short of maintaining the meaning, logic, and nuances of the original text. The translation includes partial but critical mistranslation that makes it difficult to convey the meaning of the original. Many embedded clauses, conjunctions, and sentence connectors are translated inaccurately, which results in awkward and unnatural sentence structures in the target language. Variations in the number of sentences negatively affect the readability and comprehension of the translation.
  • Score 1 The translation completely fails to convey the meaning of the entire paragraph. The translated text demonstrates a significant misinterpretation of the original text due to a critical misunderstanding of the internal logic of its complex sentences. Embedded clauses, conjunctions, and sentence connectors are inaccurately translated, resulting in unnatural and awkward sentence structures in the target language. The translated text unnecessarily alters the number of sentences so that the overall comprehensibility and accuracy of the translation is critically compromised.
C.4 Content Accuracy
Preservation of the meaning of the text, with particular attention to the proper and comprehensive translation of various literary expressions from the original text. Proper translation might involve some additions and omissions, but it should refrain from adding or omitting substantial words or phrases that constitute the content and context of the original text.
  • Score 5 The translated text fully preserves the essential meaning of the original text, demonstrating thorough comprehension of the literary text. There may be additions or omissions of particular words in the translation, but these choices are made only when they significantly enhance content accuracy in translation. On the level of meaning and literary comprehension, none of the words or phrases in the original should be missing or misinterpreted in the translation.
  • Score 4 The translated text largely preserves the essential meaning of the original text, reflecting a good overall literary comprehension. There may be some minor omissions or additions of words or expressions, but these do not significantly alter the content and context of the original text.
  • Score 3 The translated text partially preserves the essential meaning of the original text, indicating some, if not critical, miscomprehension of the original text. There are some omissions or unnecessary additions of words or expressions, which somewhat alter the content and context of the original text.
  • Score 2 The translated text barely maintains the essential meaning of the original text due to omissions or inaccurate additions of words or phrases that reflect some literary misinterpretation. The translated text misinterprets some of words or idiomatic expressions in the original text, reflecting a significant misunderstanding of the literary work.
  • Score 1 The translated text fails to convey the essential meaning of the original text due to critical omissions or inaccurate additions, revealing a fundamental lack of literary comprehension of the original text. Meaningful words or phrases in the original are either missing or replaced with expressions that do not accurately capture the content and context of the original text.
Appendix D Detailed Rubric for Step 2
This is the rubric used for whether a given translation satisfied the criteria mentioned in LLM-generated questions.
Score 3: The translation meets the criteria. Even if the translation has some issues in other aspects including naturalness, it properly conveys the meaning and affective impact of the original text, especially with regards to the aspect emphasized in the question. This score does not indicate a perfect overall translation but success in addressing the specific aspect in the given question.
Score 2: The translation has some issues relevant to the specific aspect emphasized in the question, but mostly meets the criteria. The translation may include some awkward expressions when it comes to idiomatic use of Korean language, but mostly handles the focused aspect well with occasional mistakes or lack of precision.
Score 1: The translation fails to meet the criteria due to critical errors or mistakes relevant to the given criteria. The translation misses the mark especially with regards to the specific aspect emphasized in the given question due to a limited or flawed understanding of the particular element being evaluated.
Appendix E Annotation Guidelines and User Interface
E.1 Annotation Guidelines for Step 1: Common Mistake Patterns
Notice that the human annotators have access to both this annotator’s guideline and the rubric when they do annotation.
Basic Directions and Principles
  1. 1.
    You will be asked to evaluate literary translation on a 1-5 scale.
  2. 2.
    In the annotation UI, you will be given 1) the original text, 2) translation, 3) human translation(for reference), and 4) the criteria and rubric for evaluation.
  3. 3.
    You will be asked to give separate scores for all 4 criteria per translation. The rubric for each criterion will be shown on the screen. The human translation is provided on the UI just for reference, but note that they are not always scored as 5.
  4. 4.
    Scores for each assessment criterion should be independent of assessments in other criteria. (Even if the translation as a whole is not an ideal one, you should give score 4 or 5 if the translation aligns with the description for score 4 or 5 for a specific criterion.)
Detailed Guidelines and Examples
  1. 1.
    A lexical choice score of 4 is given when there are some awkward words (up to 3) in the translation, but the original meaning can be accurately captured. A lexical choice score of 3 is awarded when the translation contains a word (up to one) that misinterprets the original meaning but does not detract from the overall meaning.
  2. 2.
    Syntax and Grammar are separate from words and are scored based on whether the sentence form or structure is appropriate to capture meaning. For example, if the words are all translated correctly, but the sentence form and structure make it difficult to understand the meaning, a low score is given. In contrast, a high score is given if the sentence structure and grammar are appropriate, even if the words are not.
  3. 3.
    If there are no sentences in the paragraph that can be evaluated using the Honorifics criteria, then the score is 5.
  4. 4.
    If the use of honorifics is a translator’s choice, it can be given high scores even if it is not the same as human translation.
  5. 5.
    Content Accuracy is linked to the broader scope of accuracy and looks mainly at whether anything is missing/added in the translation. If nothing is missing/added, then evaluate how well the translation conveys the overall flow of the original, leaving out the details.
  6. 6.
    When evaluating lexical choices, phonetically writing an English word in Korean ({CJK*}UTF8mj팬트리>pantry) does not count as an error. However, it can be counted as an error if it hinders the reading experience because it is uncommon/awkward.
  7. 7.
    The honorifics in the narration were evaluated using the syntax criteria.
E.2 Annotation Guidelines for Step 2: LLM Generated Question Answering
Similar to Step 1 annotation, annotators also had access to the rubric alongside this guideline.
Basic Directions and Principles
  1. 1.
    You will be given the original text, a human translation of the given text, and a machine-translation of the given text.
  2. 2.
    You will then be given a question that addresses a specific aspect of the translation. Those specific aspects may include: historical and cultural context, imagery and descriptive quality, characterization, interpersonal dynamics, idiomatic naturalness, subtle implications, narrative pacing and rhythm, stylistic resonance, and coherence.
  3. 3.
    After reading the original text and the machine-translation carefully, give 1 to 3 scores based on your evaluation of the translation, focusing on the criteria relevant to the given question.
  4. 4.
    When you give scores, focus only on the specific aspect mentioned in the given question when you give scores.
  5. 5.
    Refer to the rubric above for what each score means.
Specific Guidelines
  1. 1.
    Score 3 does not necessarily signify “perfect translation.” Even when the translation includes some awkward expressions in Korean, give 3 if the translation satisfies the criteria conveyed through the given question.
  2. 2.
    When the given question does not seem to apply directly to the given passage, evaluate the overall quality of translation to give 1 to 3 scores.
  3. 3.
    When the question doesn’t seem to clearly reflect the content of the translation, make proper judgments about the assumptions of the question and answer accordingly.
  4. 4.
    If the question asks you to evaluate multiple facets, and part of the criteria is met, give 2
  5. 5.
    Interpersonal dynamics–consistent and proper=3 / consistent but not accurate, or proper but partly inconsistent=2 / inconsistent and exhibits misunderstanding=1
  6. 6.
    In terms of tone, if the overall flow matches the original, give 3.
  7. 7.
    Even when the translations are awkward in terms of idiomaticity, give 3 if it meets the criteria
  8. 8.
    When the translation totally fails to capture the content of the original text, give 1.
  9. 9.
    When there are multiple verbs in the translation and the question asks about using proper verb endings, consistent and proper=3, consistent but slightly awkward=2, inconsistent and/or awkward verb endings=1
  10. 10.
    When the translation does not include enough information to answer the questions (when the question is about a broader context that is not clear in the given paragraph) > evaluate the overall quality of the translation
  11. 11.
    In case there are random foreign words or expressions included in the translation: if the foreign language parts are limited to the names or other proper nouns, give 2 or 1 depending on the quality of the translation in terms of the aspect the question focuses on. Do not give 3 to these imperfect Korean translations. if one can understand the content and context of the overall translation (even if it displays awkwardness or some mistranslation), do not mark “the translation has serious issues.” Give it 2 or 1 depending on the quality of the translation especially in terms of the aspect the question focuses on. if the entire translation includes serious grammatical errors or complete mistranslation that makes it impossible to understand the entire passage, give 1.
  12. 12.
    If a significant portion of the translation is in languages other than Korean or if the text has missing parts (i.e., the entire last line/sentence is missing) mark “This translation has serious issues” and do not give 123 scores. Just click on Submit after checking “Yes” to “serious issues” question.
  13. 13.
    If there is something wrong with the question (i.e., the question is not a question but a category or a phrase that falls short of constituting a question), do not answer that question. If only one or two of the questions from the three questions have this issue, do not give 123 scores to those questions but give 123 scores to the rest of the questions. Do not mark the entire item as “This translation has serious issues.”
  14. 14.
    If the passage does not offer enough context to answer the questions properly (primarily because the question is about the broader context of the story), give 3 (unless the translation fails to convey the content of the original text).
  15. 15.
    Although it is difficult to ignore awkward expressions or grammatical imperfections in terms of fluency in Korean, always focus on the specific aspect the question asks you to evaluate.
E.3 Annotation Platform
We use Label Studio framework Tkachenko et al. (2020-2022), hosted on Huggingface platform777huggingface.co, to create an annotation platform for all our annotations.
E.4 Annotated Samples
After the annotation, and some post-processing to discard some samples, we had 188 data points for RULER (Step 1) and 543 data points for validating VERSE (Step 2).
We will release the human annotation for future researchers.
(a) Platform for Step 1 Annotation (1 to 5 MQM)
(b) Platform for Step 2 Annotation (1 to 3 MQM)
Figure 4: Annotation platform for annotations in Section 5
Appendix F Model Name Details
Here we list the full names of the models we used in the paper: gpt-4o-2024-05-13, claude-3-5-sonnet-20240620, claude-3-5-haiku-20241022, gpt-4o-mini-2024-07-18, claude-3-haiku-20240307, gemma-2-27b-it, gemma-2-9b-it, Llama-3.1-8B-Instruct.
F.1 Baseline Details

### v3

@@ 5.2 VERSE: Can LLMs Answer Literary Questions with Subtle Nuance?
As shown in Table 4, language models are much better at correlating with human preference in understanding literary translation than the traditional MT metrics. However, there remains a gap between expert human-level agreement (Table 2). In-depth analysis of GPT-4o (the best-performing model above) performance shows that the model has a particularly lower recall than inter-annotators for Score 1 than for Score 2 and 3 (Table 5), which means GPT-4o cannot effectively spot translations that critically fail to satisfy specific criteria. However, the recall drop was not as significant as in the case of honorifics in Step 1 Honorific criteria.
Human-Human Human-Model
Label Support Prec Rec F1 Prec Rec F1
1 197/189 0.81 0.75 0.78 0.77 0.47 0.59
2 182/171 0.52 0.46 0.49 0.38 0.51 0.43
3 185/203 0.64 0.78 0.70 0.59 0.65 0.62
Avg 188/188 0.66 0.66 0.66 0.58 0.54 0.55
Table 5: Comparison of classification metrics between human annotators (Human-Human) and GPT-4o no CoT predictions (Human-Model). Support shows the average number of instances for human/model evaluations.
5.3 Ablation of Different Components in the Framework
A set of ablations tables in Appendix E shows that:
  • Chain-of-thought (CoT) does not help: Adding CoT reasoning generally fails to improve evaluation quality in either RULER or VERSE; in some cases it even slightly reduces performance. This holds true for even the ‘thinking’ reasoning models. See Table 11.
  • A clear grading rubric matters more than references: Removing the rubric hurts results much more than removing the human reference translation, showing that structured evaluation guidelines are essential (Table 12).
  • Different effects across steps: In Step 1 (RULER), human references slightly help, specially with honorifics; but in Step 2 (VERSE), models often perform better without references, suggesting different dynamics in rubric-based error detection vs. nuanced question answering (Table 12).
  • Not much gain from few-shot prompting: Providing few-shot examples does not consistently boost performance in either step, and in some cases decreases it (Table 10).
@@ 6.1 Experiment Setting
For all the experiments in this section, we use the GPT-4o as the evaluator model for all experiments in zero-shot no CoT setting. [^5] For VERSE, Claude-Sonnet 3.5 is slightly better using CoT, however, using chain of thought becomes prohibitively expensive for large scale evaluation, which is why we resort to zero shot no CoT GPT-4o, whose performance is very close. We also did not use reference human translation because in Step 2 reference-free configuration performs slightly better.
@@
To test the efficacy of our framework, we use several language models and Google Translate to translate 10 stories from our dataset. Following the best practices of Karpinska and Iyyer (2023), we prompt language models with five examples and then generate translations. We also append a GPT-4o generated summary of the story for context.
6.2 VERSE Provides Fine Grained Literary Evaluation
@@
Step 2 (VERSE) of our framework permits a novel fine-grained interpretation into literary translation, which is otherwise hard to capture through a single number or traditional machine translation metric like fluency, adequacy, etc. Figure 4 shows one such use case of such approach to quickly inspect and visualize different dimensions present in the translation. Moreover, because the language model evaluates  10 questions per paragraph, VERSE estimations are also robust.
@@ 6.3 LLMs prefers LLM generated translation
Table 6 shows the aggregated scores for each Step 1 criteria for each model. Note that since our proposed framework can be run in reference free fashion, we evaluate the reference translations as well.
The results are surprising. Overall, Claude-3.5 Sonnet (as translator) gets the highest scores, followed by GPT-4o, with almost perfect scores in several dimensions. Human translation receives lower scores than the bigger models. To better understand the cause, we sample 50 paragraphs from the story “What is Remembered”, whose human translation receive a particularly lower score than the best performing LLM translation by Claude-Sonnet 3.5. In a controlled A/B setting, one human annotator picked the gold translation as the better one 39 out of 50 times, and the Claude-Sonnet 3.5 translation was picked only 4 times.
One potential reason why evaluator LM is rating human translation lower might have to do with the nature of human translations. Human translations often deviate significantly from their original text in order to properly convey the linguistic nuances. Human inspection reveals that this is indeed the case for the “What is Remembered” story. Moreover, for all of our stories, a casual inspection reveals that human translations are clearly better than the best machine translations in majority of the cases.
@@
This suggests that while current LLMs might be good at spotting significant differences in translation qualities, they still often fail to distinguish translations where the difference is more subtle and nuanced. This can be observed from the relatively lower scores provided to Google Translate and GPT-3.5, which usually produce very poor Korean translation as seen upon human inspection.
Transl. Model Hon. Syn. Lex. Con. Step 2
Human 98.97 88.21 89.74 83.08 89.62
GPT-4o 99.49 97.44 94.36 97.95 92.41
Claude-3.5-Son. 100.00 98.97 98.97 98.97 95.55
Claude-3.5-Hai. 100.00 94.87 96.41 96.92 92.76
Claude-3-Haiku 98.46 88.72 85.64 89.23 88.11
GPT-4-Mini 99.49 90.26 88.21 92.31 87.51
GPT-3.5 91.79 57.44 56.92 60.00 67.56
Google Trans. 97.95 71.79 69.23 76.41 75.53
Table 6: Average scores (in percentage) across different models on four criteria of translation quality, computed over the 10 evaluation stories (501 paragraphs). Note that while LLMs can discriminate between very good and very bad translations (eg, GPT-4o vs GPT-3.5), its ability to distinguish between high-quality translations is limited. The rightmost column is averaged scores on Step 2 VERSE evaluation.
LLM’s tendency to overestimate scores
Translation Model Hon. Syn. Lex. Con. Style Char. Hist. Img. Comm. Ling. Pace Nuance Cons. Mean
gpt-4o 99.49 97.44 94.36 97.95 89.70 92.16 90.58 90.76 93.09 88.36 96.42 92.00 98.64 92.41
[…]
claude-3.5-sonnet 100.00 98.97 98.97 98.97 94.42 95.35 92.51 94.12 96.12 93.21 98.91 95.29 100.00 95.55
[…]
claude-3.5-haiku 100.00 94.87 96.41 96.92 90.30 92.46 90.47 91.70 93.53 87.39 97.13 92.86 99.03 92.76
[…]
claude-3-haiku 98.46 88.72 85.64 89.23 85.82 87.06 86.84 85.33 89.40 80.61 93.93 87.45 96.51 88.11
[…]
gpt-4-mini 99.49 90.26 88.21 92.31 82.79 87.15 85.91 85.73 88.06 81.82 92.02 87.37 96.71 87.51
[…]
gpt-3.5 91.79 57.44 56.92 60.00 64.48 67.03 67.54 65.68 68.06 63.27 69.28 67.29 75.39 67.56
[…]
google-translate 97.95 71.79 69.23 76.41 71.39 73.14 74.68 74.79 76.21 68.00 79.18 75.76 86.63 75.53
[…]
human 98.97 88.21 89.74 83.08 88.48 89.40 87.02 88.29 90.84 84.12 94.06 88.78 95.54 89.62
Table 7: Aggregated scores for the ten stories on RULER (Hon. - Con.) and VERSE (Style - Mean). Mean only refers to the mean of the second block, VERSE. Gray rows indicate aggregated scores when the reference human translation was also provided during evaluation. Scores for human translation were calculated in reference free manner. The model name represents the model that generated the translation. Inclusion of reference translations leads to lower scores in RULER, whereas it yields higher scores in VERSE. More interestingly, human translations are rated poorer than some machine translations.
Table 7 shows that evaluator LLM GPT-4o tends to overestimate the quality of LLM generated translations over human translations. An interesting pattern is observed with the direction of overestimation. For RULER (Step 1) scores, providing reference translation decreases the average score, while for LLM generated QA in VERSE, the average score increases. This is in line with the results observed in the ablations in Table 12 where MSE increases when reference is not provided for RULER (Step 1), and decreases for VERSE (Step 2).
Notice, however, that the relative ranks of the models do change in a few cases in these two settings, especially in the middle ranks. The best and worst-performing models’ ranks do not change.
This observation supports our previous observation in Section 6 that GPT-4o as a judge can differentiate between obviously good and obviously bad translations, however, its capability to detect differences in similarly (high) quality machine translations is not yet as robust.
@@
Although our study focuses on English–Korean, the two-step framework is adaptable to other language pairs. To provide interpretability suited specifically for literary translation, the framework introduces a novel design that combines a general, language-agnostic component with a culturally-specific rubric-based evaluation. VERSE requires little to no modification and can operate with minimal human supervision. RULER, though it operates on language-specific rubrics, can be adapted to other languages with adjustments that reflect their particular linguistic and cultural contexts.
@@ 7 Conclusion
This study evaluates the framework only on English–Korean. A central challenge for future work is to test how well LLMs can generate literary-specific questions in non-English settings and answer them in ways that respect the cultural conventions of the target language. It is possible that LLMs will not be able to address high-quality questions in low-resource languages or languages that they haven’t been trained on extensively. Similarly, it is also possible that the framework works even better in higher-resource languages. Applying our two-step framework to diverse language pairs will further reveal its capacity to capture the rich cultural variation that defines literary translation.
Appendix A Dataset Detail
The dataset detail can be found in Table 8.
Title Author Para. Count Translator(s) Orig. Trans.
[…]
@@ Appendix A Dataset Detail
[…]
@@
Table 8: The details of the stories being used in our work. The number of paragraphs and when they were published are shown. Three stories in our corpus have two different translations, for which we included the names of both translators. Gray row stories were used for human evaluation experiments in Section 5. Dark gray on top were used to create the few-shot examples for all the experiments in Section 4 and 6. The other five lighter gray works were part of the human evaluation in 4 but their paragraphs were not part of the few-shot examples. We leave the five other stories solely for the purpose of our analysis in Section 6.
Appendix B Model Name Details
Model Exact version string
GPT-4o (OpenAI, 2024) gpt-4o-2024-05-13
GPT-4.1 (OpenAI, 2025a) gpt-4.1-2025-04-14
o3 (OpenAI, 2025b) o3-2025-04-16
Claude-3.5 Sonnet (Anthropic, 2024b) claude-3-5-sonnet-20240620
Claude-3.5 Haiku (Anthropic, 2024c) claude-3-5-haiku-20241022
GPT-4o Mini (OpenAI, 2024) gpt-4o-mini-2024-07-18
Claude-3 Haiku (Anthropic, 2024a) claude-3-haiku-20240307
Gemma 2 (27B) (Google DeepMind and Google Research, 2024) gemma-2-27b-it
Gemma 2 (9B) (Google DeepMind and Google Research, 2024) gemma-2-9b-it
Llama 3.1 (8B) (Grattafiori and et al., 2024; Meta AI, 2024) Llama-3.1-8B-Instruct
DeepSeek V3 (DeepSeek-AI et al., 2024) deepseek-chat-v3
DeepSeek R1 (Guo et al., 2025) deepseek-r1
Table 9: LLMs and exact versions used in the evaluation experiments.
Table 9 shows the names and the details of the models used in our work.
B.1 Baseline Details